# Large Vision-Language Model Alignment and Misalignment: A Survey Through the Lens of Explainability

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have demonstrated remarkable capabilities in processing both visual and textual information. However, the critical challenge of alignment between visual and linguistic representations is not fully understood. This survey presents a comprehensive examination of alignment and misalignment in LVLMs through an explainability lens. We first examine the fundamentals of alignment, exploring its representational and behavioral aspects, training methodologies, and theoretical foundations. We then analyze misalignment phenomena across three semantic levels: object, attribute, and relational misalignment. Our investigation reveals that misalignment emerges from challenges at multiple levels: the data level, the model level, and the inference level. We provide a comprehensive review of existing mitigation strategies, categorizing them into parameter-frozen and parameter-tuning approaches. Finally, we outline promising future research directions, emphasizing the need for standardized evaluation protocols and in-depth explainability studies.
@@ A Survey Through the Lens of Explainability
Dong Shu1, Haiyan Zhao2, Jingyu Hu3, Weiru Liu3, Lu Cheng4, Mengnan Du2 1Northwestern University, 2New Jersey Institute of Technology, 3University of Bristol, 4University of Illinois Chicago dongshu2024@u.northwestern.edu,  mengnan.du@njit.edu
@@ 1 Introduction
Recent Large Vision-Language Models (LVLMs) have achieved significant progress in multimodal understanding. Models such as GPT-4V OpenAI , Gemini Team et al. (2023), LLaVA Liu et al. (2024b), MiniGPT-4 Zhu et al. (2023), and LLaMa 3.2 Dubey et al. (2024) demonstrate unprecedented capabilities in tasks like image captioning and visual question answering, not only processing visual and textual information independently but also reasoning across these modalities. These advances are built upon two fundamental pillars: large language models (LLMs) and vision encoders. LLMs such as GPT-3.5 Brown (2020), LLaMA Touvron et al. (2023a), LLaMA 2 Touvron et al. (2023b), Vicuna Chiang et al. (2023), and Qwen Bai et al. (2023) revolutionized natural language processing, while vision encoders like CLIP Radford et al. (2021) have transformed the ability to create aligned visual and textual representations, enabling effective vision-language understanding.
The key challenge in developing effective LVLMs lies in achieving proper alignment between visual and linguistic representations Liu et al. (2024a). The predominant approach involves using representation alignment techniques, where visual features from an image encoder and textual representations from an LLM are mapped into a shared embedding space, typically matching the LLM’s embedding dimensions Jia et al. (2021); Yang et al. (2022); Shu et al. (2024). Once both modalities are mapped into this shared space, alignment can be achieved through various training objectives and architectural designs that encourage the model to understand and reason about cross-modal relationships. This method has gained popularity due to its straightforward approach and generalizability across different model architectures.
@@
In this survey, we present a structured framework for understanding and addressing alignment challenges in LVLMs from an explainability perspective. We first examine the fundamentals of alignment, including its representational and behavioral aspects, training procedures, and theoretical foundations. We then analyze misalignment phenomena across three semantic levels: object, attribute, and relational misalignment. Our investigation reveals that misalignment stems from challenges at the data level (e.g., quality and balance issues), model level (e.g., architectural limitations and ability gaps), and inference level (e.g., task discrepancies). We review existing mitigation strategies and outline future research directions, emphasizing the need for standardized evaluation protocols and in-depth explainability studies.
Figure 1: Overview of the three-stage LVLM training process, showing the progression from contrastive learning of visual-text encoders, through adapter fine-tuning with frozen components, to end-to-end model training.
@@ 2 Alignment of LVLMs
In this section, we examine alignment in LVLMs across four essential dimensions. First, we define the concept of alignment in LVLMs. Second, we detail the procedural stages through which alignment is achieved in practice. Third, we explore the theoretical foundations that make alignment possible between visual and linguistic modalities. Finally, we discuss methods for measuring and evaluating alignment in LVLMs.
@@ 2.1 What is Alignment?
  • Representational alignment refers to the degree of correspondence between visual representations $v\in\mathcal{V}$ and textual representations $t\in\mathcal{T}$ within the model’s internal embedding space $\mathcal{E}$. When well-aligned, the visual features extracted from an image and the textual embeddings of its corresponding description occupy nearby regions in the shared latent space, exhibiting high semantic similarity $d(v,t)$ where $d$ is a similarity metric. This internal alignment enables the model to establish meaningful connections between visual and linguistic information at a fundamental level.
@@ 2.2 How is Alignment Achieved?
The development of alignment in LVLMs progresses through three major stages (see Figure 1), each is built upon its predecessor to achieve increasingly sophisticated cross-modal integration.
@@
Stage 3: End-to-End Fine-tuning.  The final stage involves comprehensive fine-tuning of the entire system, including the visual encoder, adapter, and LLM components together. This comprehensive approach allows for deeper integration and more sophisticated alignment between all components. It enables the model to learn task-specific optimizations that require coordinated adjustments across all modules. Through this process, the model develops more advanced cross-modal understanding capabilities and facilitates the emergence of emergent behaviors that arise from the deep integration of visual and linguistic processing. This stage often results in the highest performance but requires careful balancing to avoid catastrophic forgetting or degradation of pre-existing capabilities.
@@ 2.3 Why is Alignment Possible?
Theoretical Perspective.  From a theoretical standpoint, visual and textual data are different projections of the same underlying reality. As Huh et al. argue in their Platonic Representation Hypothesis Huh et al. (2024), all modalities are measurements of a real world that generates our observations. When humans create images or write text, they are encoding information about this same reality, albeit through different measurement processes. Although these modalities appear distinct on the surface, they fundamentally capture overlapping semantic information about the same world state. This shared origin in physical reality, combined with the fact that humans generate both types of data to describe their observations of the world, provides the theoretical foundation for why these modalities can be meaningfully aligned in a common representation space.
@@
2.4 How to Measure Alignment?
Behavioral Level.  The behavioral level measures alignment through the model’s performance on various downstream tasks and benchmarks, using both direct comparisons and automated evaluation systems. The strength of alignment directly impacts the LVLM’s performance, as better alignment typically leads to improved task outcomes. These measurements generally involve comparing the model’s outputs against ground truth labels, either through direct comparison or using evaluation models to simulate human judgment. Numerous benchmarks have been developed to assess LVLM alignment across a range of tasks, from coarse-grained evaluations (e.g., object existence) to fine-grained assessments (e.g., color, count, spatial relations). Examples of such benchmarks include POPE Li et al. (2023b), CHAIR Rohrbach et al. (2018), MME Fu et al. (2023), MMHal-Bench Sun et al. (2023), and LLaVa-Bench Liu et al. (2024b). In addition to traditional benchmarks, advanced evaluation models like GAVIR Liu et al. (2023), CCEval Zhai et al. (2023) and HaELM Wang et al. (2023) provide sophisticated assessments by considering context and evaluating responses comprehensively, similar to human evaluators. The flexibility and diversity of evaluation models enable thorough measurement capabilities needed for open-ended questions.
@@ 3.2.1 Dataset Level
Data quality and distribution patterns play crucial roles in contributing to misalignment between visual and language representations in LVLMs. Several key dataset factors can impede the model’s ability to form accurate associations between visual inputs and textual descriptions, affecting both training effectiveness and inference performance.
@@
  • Data False Negative: False negatives in the dataset further complicate alignment, as negative image-text pairs, though not perfectly matching, share overlapping components Liang et al. (2022); Byun et al. (2024). During training, embeddings of positive pairs are drawn closer together, while those of negative pairs are pushed apart. This binary method can suppress latent similarities within false negatives, reducing the model’s capacity to generalize and effectively align diverse modalities.
@@ 3.2.2 Model Level
  • Separate Training: Prior to being integrated as a single LVLM, the visual encoder and the LLM are typically pre-trained separately on distinct single-modality datasets. While this approach offers advantages such as efficiency and modularity, it leads to each model developing its own biased representations and understanding of the world, shaped by its respective single-modality data Bordes et al. (2024).
@@
  • Knowledge Conflict: A significant challenge arises from knowledge conflicts between the visual encoder and language model components of LVLMs. These conflicts emerge when the visual encoder’s direct perception of image content contradicts the prior knowledge embedded in the LLM’s parameters during pre-training Zhu et al. (2024b); Ghosh et al. (2024). For example, when an image contains a green tomato, the visual encoder accurately detects its color, but the LLM may resist this information since it has been predominantly trained on texts describing ripe, red tomatoes. This misalignment between observed visual evidence and learned textual priors can manifest in various ways: the model might incorrectly describe the tomato as red despite clear visual evidence, generate hesitant or self-contradicting descriptions, or attempt to rationalize the discrepancy by making unwarranted assumptions about the tomato’s ripeness stage.
3.2.3 Inference Level
@@
 Woo et al. (2024a), Kim et al. (2024b), VDGD Ghosh et al. (2024) , leaf2, text width=42em] ] ] [Inference-Based [e.g., VTI Liu et al. (2024c), ICT Chen et al. (2024a), MetaToken Fieback et al. (2024) , leaf2, text width=49.6em] ] [Decoding-Based [e.g., AVC Woo et al. (2024b), DAMRO Gong et al. (2024), VCD Leng et al. (2024), ICD Wang et al. (2024d),
@@ 4 Mitigation Methods
Building upon our analysis of misalignment causes in LVLMs, we now examine strategies for mitigating these challenges (see Figure 3). These mitigation approaches can be categorized into two groups: parameter-tuning alignment methods and parameter-frozen alignment methods. Parameter-tuning alignment involves modifying specific components within the LVLM architecture to reduce misalignment through targeted parameter updates. In contrast, parameter-frozen alignment methods address misalignment while maintaining the LVLM’s original parameters unchanged, offering solutions that preserve the model’s structure while improving its cross-modal alignment capabilities.
@@ 4.2 Parameter Frozen Alignment
Augment-based Methods.  As analyzed in Section 3, insufficient input of image information is one of the primary causes of misalignment, leading to poor visual understanding. To address this, retrieval-augmented generation (RAG) methods have been adapted to dynamically integrate external knowledge into LVLMs through retrieved results Qu et al. (2024); Chen et al. (2023a); Ramos et al. (2023a, b); Sarto et al. (2024); Yang et al. (2024). By reranking the similarity of image-text pairs, RAG approaches provide more visual context and guidance to the model. Similarly, other methods rely on generating approach to enrich the input with additional information. For instance, Zhao et al. (2024d), Li et al. (2024a), and Zhao et al. (2023a) propose integrating an auxiliary model to generate relevant information based on the image. Alternatively, methods such as RITUAL Woo et al. (2024a) bypass the need for external models. It enhances the model’s exposure to diverse visual contexts by applying random transformations to input images. Additionally, approaches like Kim et al. (2024b); Ghosh et al. (2024) employ self-generated textual descriptions appended to the input prompt, ensuring the model has sufficient knowledge to answer questions accurately without generating plausible but incorrect responses.
@@
Inference-based Methods.  Some methods operate in the model’s latent space during the inference process, prior to decoding, by intervening in both visual and textual representations to improve alignment. For instance, Visual and Textual Intervention (VTI) Liu et al. (2024c) pre-computes intervention directions using a small set of examples and applies them during inference to enhance feature stability and vision-text alignment, without requiring additional training. Similarly, Image-Object Cross-Level Trusted Intervention (ICT) Chen et al. (2024a) introduces a lightweight mechanism that intervenes in the model’s attention at both image and object levels, applying targeted activation shifts to selected attention heads. Since they operate directly on the model’s internal representations, they can make precise adjustments to improve alignment without disrupting the model’s broader language understanding capabilities. This makes inference-based methods particularly effective at reducing misalignment while preserving the model’s ability to generate fluent and contextually appropriate responses.
@@
Decoding-based Methods.  Another widely used approach for mitigating misalignment involves modifying decoding process. These methods often target issues of imbalanced attention. However, the imbalance attention between what still remain debated. Some researchers argue that the model over-focuses on irrelevant image tokens, such as background elements or unimportant details Woo et al. (2024b); Gong et al. (2024). However, the prevailing view is that the model prioritizes textual tokens over visual ones, neglecting critical visual information Leng et al. (2024); Wang et al. (2024d); Kim et al. (2024a); Liu et al. . Despite these differences in interpretation, most decoding-based methods employ similar contrastive decoding techniques to rebalance attention between modalities, typically by reducing attention to textual tokens while enhancing focus on visual tokens. This approach, however, contrasts with inference-based methods, which avoid reducing attention to textual information and instead preserve the model’s overall language understanding. Another interesting observation is that, while decoding-based methods typically lead to similar approaches, in some cases, they can result in fundamentally divergent strategies. For instance, OPERA Huang et al. (2024a) hypothesizes that the model over-relies on summary tokens, instead of focusing visual tokens. However, text summarization is SGD’s solution Min et al. (2024) to mitigate misalignment. It uses summarization to shorten textual context and helps model shift focus toward visual information. This divergence underscores how subtle differences in understanding the root causes of misalignment can result in contradicted methodologies.
@@
Post-decoding Methods.  Lastly, post-decoding approaches present a broader range of hypotheses about the causes of misalignment, tackling issues ranging from data-level biases to model-level deficiencies. Methods such as LURE Zhou et al. (2023b) and Woodpecker Yin et al. (2023) exemplify this category. LURE focuses on addressing object hallucinations by revising the generated text, identifying hallucinatory content, and reconstructing less biased outputs. Woodpecker employs a five stages validation mechanism to extract and correct inconsistencies in the generated response. Despite their specific details, these methods converge on a shared strategy, which involves modifying the model’s outputs after decoding without altering the its parameters or architecture, making them easily adaptable to various LVLMs. This flexibility lies in their goal-oriented nature, as they directly target specific misalignment phenomena. However, this goal-oriented focus introduces a significant limitation. While it can enhance output quality, the underlying model deficiencies are left unchanged, restricting generalization and limiting performance on tasks beyond post-decoding corrections.
5 Future Research Directions

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) have demonstrated remarkable capabilities in processing both visual and textual information. However, the critical challenge of alignment between visual and textual representations is not fully understood. This survey presents a comprehensive examination of alignment and misalignment in LVLMs through an explainability lens. We first examine the fundamentals of alignment, exploring its representational and behavioral aspects, training methodologies, and theoretical foundations. We then analyze misalignment phenomena across three semantic levels: object, attribute, and relational misalignment. Our investigation reveals that misalignment emerges from challenges at multiple levels: the data level, the model level, and the inference level. We provide a comprehensive review of existing mitigation strategies, categorizing them into parameter-frozen and parameter-tuning approaches. Finally, we outline promising future research directions, emphasizing the need for standardized evaluation protocols and in-depth explainability studies.
@@ A Survey Through the Lens of Explainability
Dong Shu1, Haiyan Zhao2, Jingyu Hu3, Weiru Liu3, Ali Payani4, Lu Cheng5, Mengnan Du2 1Northwestern University, 2New Jersey Institute of Technology, 3University of Bristol, 4Cisco Research,5University of Illinois Chicago dongshu2024@u.northwestern.edu,  mengnan.du@njit.edu
@@ 1 Introduction
Recent Large Vision-Language Models (LVLMs) have achieved significant progress in multimodal understanding. Models such as GPT-4V OpenAI , Gemini Team et al. (2023), LLaVA Liu et al. (2024b), Claude-3.5-Sonnet Anthropic (2024), Qwen2-VL Wang et al. (2024d), and LLaMa 3.2 Dubey et al. (2024) demonstrate remarkable capabilities in tasks like image captioning and visual question answering, not only processing visual and textual information independently but also reasoning across these modalities. These advances are built upon two two key foundations: large language models (LLMs) and vision encoders. LLMs such as GPT-3.5 Brown (2020), LLaMA Touvron et al. (2023a), LLaMA 2 Touvron et al. (2023b), and Qwen Bai et al. (2023a) transformed natural language processing, while visual understanding models like Vision Transformer (ViT) Dosovitskiy et al. (2021) and vision-language models like CLIP Radford et al. (2021) have advanced the ability to process visual information and create aligned visual-textual representations respectively.
The key challenge in developing effective LVLMs lies in achieving proper alignment between visual and textual representations Liu et al. (2024a). The predominant approach involves using representation alignment techniques, where visual features from an image encoder and textual representations from an LLM are mapped into a shared embedding space, typically matching the LLM’s embedding dimensions Jia et al. (2021); Yang et al. (2022); Shu et al. (2024). Once both modalities are mapped into this shared space, alignment can be achieved through various training objectives and architectural designs that encourage the model to understand and reason about cross-modal relationships. This method has gained popularity due to its straightforward approach and generalizability across different model architectures.
@@
In this survey, we present a structured framework for understanding and addressing alignment challenges in LVLMs from an explainability perspective. We first examine the fundamentals of alignment, including its representational and behavioral aspects, training procedures, and theoretical foundations. We then analyze misalignment phenomena across three semantic levels: object, attribute, and relational misalignment. Our investigation reveals that misalignment stems from challenges at the data level (e.g., quality and balance issues), model level (e.g., architectural limitations and ability gaps), and inference level (e.g., task discrepancies). We review existing mitigation strategies and outline future directions, emphasizing the need for standardized evaluation protocols and in-depth explainability studies.
@@ 2 Alignment of LVLMs
In this section, we examine alignment in LVLMs across four essential dimensions. First, we define the concept of alignment in LVLMs. Second, we detail the procedural stages through which alignment is achieved in practice. Third, we explore the theoretical foundations that make alignment possible between visual and textual modalities. Finally, we discuss methods for measuring and evaluating alignment in LVLMs in Appendix B.
@@ 2.1 What is Alignment?
  • Representational alignment refers to the degree of correspondence between visual representations $v\in\mathcal{V}$ and textual representations $t\in\mathcal{T}$ within the model’s internal embedding space $\mathcal{E}$. When well-aligned, the visual features extracted from an image and the textual embeddings of its corresponding description occupy nearby regions in the shared latent space, exhibiting high semantic similarity $d(v,t)$ where $d$ is a similarity metric. This internal alignment enables the model to establish meaningful connections between visual and textual information at a fundamental level.
@@ 2.2 How is Alignment Achieved?
The development of alignment in LVLMs progresses through three major stages (see Figure 3 in Appendix), each is built upon its predecessor to achieve increasingly sophisticated cross-modal integration. Additionally, we present prominent real-world LVLM architectures in Appendix A.
@@
Stage 3: End-to-End Fine-tuning.  The final stage involves comprehensive fine-tuning of the entire system, including the visual encoder, adapter, and LLM components together. This comprehensive approach allows for deeper integration and more sophisticated alignment between all components. It enables the model to learn task-specific optimizations that require coordinated adjustments across all modules. Through this process, the model develops more advanced cross-modal understanding capabilities and facilitates the emergence of emergent behaviors that arise from the deep integration of visual and textual processing. This stage often results in the highest performance but requires careful balancing to avoid catastrophic forgetting or degradation of pre-existing capabilities.
@@ 2.3 Why is Alignment Possible?
Theoretical Perspective.  From a theoretical standpoint, visual and textual data are different projections of the same underlying reality. As Huh et al. argue in their Platonic Representation Hypothesis Huh et al. (2024), all modalities are measurements of a real world that generates our observations. When humans create images or write text, they are encoding information about this same reality, though through different measurement processes. Although these modalities appear distinct on the surface, they fundamentally capture overlapping semantic information about the same world state. This shared origin in physical reality, combined with the fact that humans generate both types of data to describe their observations of the world, provides the theoretical foundation for why these modalities can be meaningfully aligned in a common representation space.
@@ 3.1 Definition of Misalignment
To better illustrate real-life examples of misalignment, we have selected four prominent LVLMs and provided their generated responses in Appendix D.
@@ 3.2.1 Dataset Level
Data quality and distribution patterns play crucial roles in contributing to misalignment between visual and language representations in LVLMs. Several key dataset factors can impede the model’s ability to form accurate associations between visual inputs and textual descriptions, affecting both training effectiveness and inference performance. We list one factor below, for additional factors please refer to Appendix C.
@@ 3.2.2 Model Level
  • Knowledge Conflict: A significant challenge arises from knowledge conflicts between the visual knowledge of image and parametric knowledge of LLM. These conflicts emerge when the visual encoder’s direct perception of image content contradicts the prior knowledge embedded in the LLM’s parameters during pre-training Zhu et al. (2024b); Ghosh et al. (2024). For example, when an image contains a green tomato, the visual encoder accurately detects its color, but the LLM may resist this information since it has been predominantly trained on texts describing ripe, red tomatoes. This misalignment between observed visual evidence and learned textual priors can manifest in various ways: the model might incorrectly describe the tomato as red despite clear visual evidence, generate hesitant or self-contradicting descriptions, or attempt to rationalize the discrepancy by making unwarranted assumptions about the tomato’s ripeness stage.
@@
 Woo et al. (2024a), Kim et al. (2024b), VDGD Ghosh et al. (2024) , leaf2, text width=42em] ] ] [Inference-Based [e.g., VTI Liu et al. (2024c), ICT Chen et al. (2024a), MetaToken Fieback et al. (2024) , leaf2, text width=49.6em] ] [Decoding-Based [e.g., AVC Woo et al. (2024b), DAMRO Gong et al. (2024), VCD Leng et al. (2024), ICD Wang et al. (2024e),
@@ 4 Mitigation Methods
Building upon our analysis of misalignment causes in LVLMs, we now examine strategies for mitigating these challenges (see Figure 2). These mitigation approaches can be categorized into two groups: parameter-tuning alignment methods and parameter-frozen alignment methods. Parameter-tuning alignment involves modifying specific components within the LVLM architecture to reduce misalignment through targeted parameter updates. In contrast, parameter-frozen alignment methods address misalignment while maintaining the LVLM’s original parameters unchanged, offering solutions that preserve the model’s structure while improving its cross-modal alignment capabilities. We have conducted experiments comparing the performance and computational cost of several mitigation methods (see Appendix E for details).
@@ 4.2 Parameter Frozen Alignment
Augment-based Methods.  As analyzed in Section 3, insufficient input of image information is one of the primary causes of misalignment, leading to poor visual understanding. To address this, retrieval-augmented generation (RAG) methods have been adapted to dynamically integrate external knowledge into LVLMs through retrieved results Qu et al. (2024); Chen et al. (2023a); Ramos et al. (2023a, b); Sarto et al. (2024); Yang et al. (2024). By reranking the similarity of image-text pairs, RAG approaches provide more visual context and guidance to the model. Similarly, other methods rely on generating approach to enrich the input with additional information. For instance, Zhao et al. (2024d), Li et al. (2024a), and Zhao et al. (2023a) propose integrating an auxiliary model to generate relevant information based on the image. Alternatively, methods such as RITUAL Woo et al. (2024a) bypass the need for external models. It enhances the model’s exposure to diverse visual contexts by applying random transformations to input images. Additionally, approaches like Kim et al. (2024b); Ghosh et al. (2024) employ self-generated textual descriptions appended to the input prompt, ensuring the model has sufficient knowledge to answer questions accurately.
@@
Inference-based Methods.  Some methods operate in the model’s latent space during the inference process, prior to decoding, by intervening in both visual and textual representations to improve alignment. For instance, Visual and Textual Intervention (VTI) Liu et al. (2024c) pre-computes intervention directions using a small set of examples and applies them during inference to enhance feature stability and vision-text alignment, without requiring additional training. Similarly, Image-Object Cross-Level Trusted Intervention (ICT) Chen et al. (2024a) introduces a lightweight mechanism that intervenes in the model’s attention at both image and object levels, applying targeted activation shifts to selected attention heads. Since they operate directly on the model’s internal representations, they can make precise adjustments to improve alignment without disrupting the model’s broader language understanding capabilities. This makes inference-based methods effective at reducing misalignment while preserving the model’s ability to generate contextually appropriate responses.
@@
Decoding-based Methods.  Another common approach to mitigating misalignment involves modifying the decoding process. These methods often target issues of imbalanced attention. However, the imbalance attention between what still remain debated. Some researchers argue that the model over-focuses on irrelevant image tokens, such as background elements or unimportant details Woo et al. (2024b); Gong et al. (2024). However, the prevailing view is that the model prioritizes textual tokens over visual ones, neglecting critical visual information Leng et al. (2024); Wang et al. (2024e); Kim et al. (2024a); Liu et al. . Despite these differences in interpretation, most decoding-based methods use contrastive decoding to rebalance attention between modalities, typically by reducing attention to textual tokens while enhancing focus on visual tokens. This approach, however, contrasts with inference-based methods, which avoid reducing attention to textual information and instead preserve the model’s overall language understanding. Another interesting observation is that, while decoding-based methods typically lead to similar approaches, they can sometimes result in divergent strategies. For instance, OPERA Huang et al. (2024a) hypothesizes that the model over-relies on summary tokens, instead of focusing visual tokens. However, text summarization is SGD’s solution Min et al. (2024) to misalignment. It uses summarization to shorten textual context and helps model shift focus toward visual information. This divergence underscores how subtle differences in understanding misalignment’s root causes can lead to contradicted methodologies.
@@
Post-decoding Methods.  Lastly, post-decoding approaches present broader hypotheses about misalignment causes, tackling issues ranging from data-level biases to model-level deficiencies. Methods such as LURE Zhou et al. (2023b) and Woodpecker Yin et al. (2023) exemplify this category. LURE focuses on addressing object hallucinations by revising the generated text, identifying hallucinatory content, and reconstructing less biased outputs. Woodpecker employs a five stages validation mechanism to extract and correct inconsistencies in the generated response. Despite their specific details, these methods converge on a shared strategy, which involves modifying the model’s outputs after decoding without altering parameters or architecture, making them easily adaptable to various LVLMs. This flexibility lies in their goal-oriented nature, as they directly target specific misalignment phenomena.
5 Conclusions
In this paper, we systematically survey alignment and misalignment in LVLMs through an explainability lens. We show that achieving proper alignment involves complex interactions between data quality, model architecture, and inference procedures. We categorize misalignment into object, attribute, and relational levels, providing a clear framework to understand these challenges and develop targeted solutions. Our review of mitigation strategies highlights approaches ranging from computationally intensive parameter-tuning methods to practical parameter-frozen solutions, each with trade-offs in effectiveness and feasibility. Lastly, we outline key future directions for building robust vision-language systems (see Appendix F), including development of standardized benchmarks, deeper investigation into explainability approaches to reveal internal alignment mechanisms, and architectural innovations to address fundamental limitations in current designs.
Limitations
While this paper provides a comprehensive survey of alignment and misalignment in LVLMs, we acknowledged there are several limitations. Currently, the scope of this survey is limited to the alignment between vision and language modalities. However, real-world applications often require the integration of multiple modalities, such as audio, video, or sensor data. We plan to expand the scope to explore the challenges and alignment techniques in such complex multi-modal settings, including investigating alignment mechanisms in architectures that incorporate three or more modalities, developing new evaluation metrics for multi-modal alignment, and studying the interactions between different modalities in integrated systems.
Appendix A Overview of LVLM Architectures
Table 1 summarizes the architectural components of six prominent LVLMs discussed throughout this survey paper on model alignment and misalignment. The comparison reveals common patterns in LVLM design: transformer-based vision encoders (predominantly ViT variants), adapter modules of varying complexity to connect vision and language components, and state-of-the-art language models.
Appendix B How to Measure Alignment?
Behavioral Level.  The behavioral level measures alignment through the model’s performance on various downstream tasks and benchmarks, using both direct comparisons and automated evaluation systems. The strength of alignment directly impacts the LVLM’s performance, as better alignment typically leads to improved task outcomes. These measurements generally involve comparing the model’s outputs against ground truth labels, either through direct comparison or using evaluation models to simulate human judgment. Numerous benchmarks have been developed to assess LVLM alignment across a range of tasks, from coarse-grained evaluations (e.g., object existence) to fine-grained assessments (e.g., color, count, spatial relations). Examples of such benchmarks include POPE Li et al. (2023b), CHAIR Rohrbach et al. (2018), MME Fu et al. (2023), MMHal-Bench Sun et al. (2023), and LLaVa-Bench Liu et al. (2024b). In addition to traditional benchmarks, advanced evaluation models like GAVIE Liu et al. (2023), CCEval Zhai et al. (2023) and HaELM Wang et al. (2023) provide sophisticated assessments by considering context and evaluating responses comprehensively, similar to human evaluators. The flexibility and diversity of evaluation models enable thorough measurement capabilities needed for open-ended questions.
In Table 2, we compare several existing benchmarks. While most of these benchmarks focus on evaluating misalignment, some also assess reasoning capabilities, such as whether the LVLM can reason based on image information or whether the LVLM fully follow user input instructions. Future benchmarks should also adopt this comprehensive approach, evaluating not only misalignment but also whether mitigating misalignment impacts the LVLM’s reasoning ability, as our goal is to develop LVLMs with minimal misalignment while preserving their reasoning capabilities.
Appendix C More Reasons for Misalignment
Beyond the dataset-level issues discussed in Section 3.2.1, we present additional dataset-related challenges that may contribute to the misalignment of LVLMs.
  • Data False Negative: False negatives in the dataset further complicate alignment, as negative image-text pairs, though not perfectly matching, share overlapping components Liang et al. (2022); Byun et al. (2024). During training, embeddings of positive pairs are drawn closer together, while those of negative pairs are pushed apart. This binary method can suppress latent similarities within false negatives, reducing the model’s capacity to effectively align diverse modalities.
Table 2: Comparison of Misalignment Benchmark across Different Evaluation Dimensions. The evaluation dimensions are categorized into two main aspects: (1) Evaluating misalignment, which examines object, attribute, and relational misalignments in model outputs; and (2) Evaluating reasoning, which assesses models’ cognitive reasoning abilities and instruction following capabilities. The rightmost column indicates whether the benchmark uses traditional evaluation methods or third-model evaluation approaches.
Evaluating Misalignment Evaluating Reasoning
Object Misalignment Attribute Misalignment Relational Misalignment Cognition Reasoning Instruction Following Traditional/ Third-model
POPE Li et al. (2023b) Traditional
CHAIR Rohrbach et al. (2018) Traditional
MME Fu et al. (2023) Traditional
MMHal-Bench Sun et al. (2023) Traditional
LLaVa-Bench Liu et al. (2024b) Traditional
LVLM-eHub Xu et al. (2024) Traditional
GAVIE Liu et al. (2023) Third-model
CCEval Zhai et al. (2023) Third-model
HaELM Wang et al. (2023) Third-model
Appendix D Examples of LVLM Descriptions
Table 3: Comparison of Mitigation Methods on the POPE Benchmark. The table demonstrates the performance of the baseline model and different mitigation methods that address baseline misalignment. Performance is evaluated using Accuracy, Precision, Recall, and F1 score (with the highest score in each category underlined). Additionally, the computational cost (in seconds per input) for both the baseline and each mitigation method is provided.
POPE-Random POPE-Adversarial POPE-Popular Comp. Cost
Method Acc Precis Recall F1 Acc Precis Recall F1 Acc Precis Recall F1 Sec/Input
Baseline
LLaVa-V1.5-7B 88.4 87.5 89.7 88.6 77.9 72.6 89.7 80.3 84.8 81.7 89.7 85.5 1.27
Mitigations
SoM-LLaVA Yan et al. (2024) 89.6 89.1 90.2 89.6 81.0 76.2 90.2 82.6 84.8 81.4 90.2 85.6 1.62
SID Huo et al. (2024) 89.4 92.2 86.1 89.0 80.4 77.4 85.9 81.4 85.9 85.9 85.9 85.9 1.68
LogicCheckGPT Wu et al. (2024a) 88.0 98.3 77.3 86.5 85.0 94.1 74.7 83.3 86.7 95.1 77.3 85.3 51.5
In this section, we present examples from four leading LVLMs and illustrate their generated descriptions for given images. We have selected one closed-source model, ChatGPT-4o, accessed through its official website https://chatgpt.com/, and three open-source models, Qwen2-VL-72B-Instruct, DeepSeek-VL2, and LLaVa-1.5-13B-hf, loaded directly from Hugging Face. All implementation settings, including parameters such as temperature, were kept at their default values to ensure consistency. The experiments were conducted using an A100 PCIE 80GB GPU. To enhance computational efficiency and performance, the models were loaded in float16 precision. During the evaluation phase, each model was provided with an image along with the following prompt: ‘Describe this image in detail. Include specific objects, their attributes, and relationships.’ The input images and the corresponding descriptions generated by each model are presented in Figure 4, 5, and 6.
Appendix E Mitigation Methods Comparison
As shown in Table 3, we selected LLaVa-V1.5-7B as the baseline model and evaluated it on POPE benchmarks. We then tested three mitigation methods: SoM-LLaVA Yan et al. (2024), SID Huo et al. (2024), and LogicCheckGPT Wu et al. (2024a), to assess their effectiveness and efficiency in mitigating misalignment in the baseline model. The baseline model was loaded directly from Hugging Face, with parameters in float16 precision and all other implementation details set to default. Each mitigation method was implemented strictly according to its official GitHub repository, without modifications to core components, except for necessary adjustments such as path configurations.
We evaluated model performance using Accuracy, Precision, Recall, and F1-score, with the highest values underlined in the table. The results indicate that mitigation methods generally improve performance across most metrics compared to the baseline, particularly SoM-LLaVA and LogicCheckGPT, which achieved five and six highest scores, respectively. However, a deeper analysis reveals that some methods negatively impacted certain metrics. For instance, while LogicCheckGPT achieved the highest Precision on all benchmarks, it exhibited relatively lower Recall and F1 scores, sometimes even worse than the baseline. This occurs because LogicCheckGPT is conservative in predicting positives, meaning it only classifies a sample as positive when it is highly confident, resulting in fewer false positives but more false negatives. Interestingly, SoM-LLaVA consistently achieved the highest Recall across all benchmarks, indicating that it is less strict in determining positive cases compared to LogicCheckGPT. This highlights how different mitigation strategies lead to different prediction behaviors, emphasizing the need for future works to report performance across multiple metrics rather than focusing solely on a single high-performing metric.
Beyond performance, we also evaluated the computational cost of each mitigation method in seconds per input. While SoM-LLaVA and SID had only a minor increase in inference time compared to the baseline, LogicCheckGPT required significantly more time. This is because LogicCheckGPT relies on an external third model, such as GPT-3.5, to assess logical consistency and mitigate object misalignment, making its computational cost highly dependent on the third model’s complexity. This trade-off between performance and computational efficiency is critical, as high computational time may hinder real-world deployment of mitigation techniques. Future research should aim to balance effectiveness with efficiency to ensure practical usability.
Appendix F Future Research Directions